Circular No. 8176

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003gw IN UGC 3252 Further to IAUC 8173, M. Armstrong reports his discovery of an ap-

parent supernova on unfiltered CCD images taken on Aug. 7.114 (mag 17.1) and 7.945 UT (mag 17.3). SN 2003gw is located at α = 5^{h}11^{m}47^{s}.68, δ = +67^{o}29′22′′.2, which is ≈ 14′′ east and 6′′ north of the center of UGC 3252. Nothing is visible at this position on his images from 2000 Nov. 3 and 2003 Mar. 22 (limiting mag 19.0) or on Palomar Sky Survey plates.

SUPERNOVA 2003go Further to IAUC 8173, M. Hamuy, N. Morrell, J. Arias, and R. Barba

report that a spectrum (range 390-900 nm) of SN 2003go (cf. IAUC 8168), obtained on July 30.30 UT with the Baade telescope, shows it to be a type-IIn supernova with a blue continuum and strong Balmer lines in emission, at a redshift of 0.051. Superimposed, narrow, absorption-line components can be seen blueshifted by ∼ 600 km/s from the emission line cores of the Balmer lines. A weak He I 587.6-nm emission is also present.

V4743 SAGITTARII R. M. Wagner, LBT Observatory; S. G. Starrfield and J. Rohrbach,

Arizona State University; H. Bond, Space Telescope Science Institute; A. Karam, University of Arizona; B. Monard, Bronberg Observatory; G. Bolt, Craigie, Australia; and R. Ishioka, M. Uemura, and T. Kato, Kyoto University, report: “A comparison of optical photometry, obtained contemporaneously with Chandra x-ray observations in April and July, reveals a substantial change in the optical light curve of this classical nova and super-soft-x-ray source. Unfiltered photometry obtained on July 18, 19, and 21 and B-band photometry obtained at Cerro Tololo (1.3-m reflector + AndiCam) on July 19 show a sinusoidal modulation of amplitude 0.2 mag with a period of 0.281 ± 0.003 day, which is likely to be the orbital period. Minimum light corresponds to JD 2452839.321. Another variation with a period of 69 ± 3 min is superposed on the orbital modulation. Photometry obtained at Cerro Tololo with the same equipment on seven nights between Apr. 5.3 and 20.4 UT shows erratic variations that are not periodic or consistent with the July results. No statistically significant variation is detected near the x-ray period of 1325 s (Ness et

al. 2003, http://xxx.lanl.gov/abs/astro-ph/0308017) or at twice this

period. The recent photometric behavior is similar to some intermediate polar cataclysmic variables. Additional observations are encouraged.”

2003 August 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green